# Notes on heating phase dynamics in Floquet CFTs and Modular quantization

### v1

@@ Abstract
In this article, we explore the connection between the heating phase of periodically driven CFTs and the Modular Hamiltonian of a subregion in the vacuum state. We show that the heating phase Hamiltonian corresponds to the Modular Hamiltonian, with the fixed points mapping to the endpoints of the subregion. In the bulk dual, we find that these fixed points correspond to the Ryu-Takayanagi surface of the AdS-Rindler wedge. Consequently, the entanglement entropy associated to the boundary interval within two fixed points exactly matches with the Bekenstein-Hawking entropy of AdS-Rindler. We observe the emergent Virasoro algebra in the boundary quantization of the Modular Hamiltonian has a striking similarity with the emerging near Horizon Virasoro algebra. This is a consequence of the fact that while obtaining the boundary Virasoro algebra, a cut-off with conformal boundary condition around the fixed point is introduced, which in the bulk is related to a stretched horizon, with an emergent two-dimensional conformal symmetry. We also argue that as one tunes the parameter space of Floquet Hamiltonians to transition from the non-heating to the heating phase the operator algebra type changes from Von Neumann type $I$ to $III_{1}$ factor, providing a non-equilibrium analogue of the Hawking-Page transition.
@@ 1 Introduction
The study of Floquet CFTs has recently garnered considerable attention as analytically tractable models of physics out of equilibrium. At stroboscopic times, the evolution of such systems is governed by an effective Hamiltonian, referred to in the literature as the Floquet Hamiltonian $H_{F}$. One of the novel features of their dynamics is the existence of three different dynamical phases, characterized by qualitatively different temporal growths of the entanglement entropy and energy. In particular, one of the phases, aptly referred to as the heating phase, is characterized by an exponentially growing energy and linearly growing entanglement entropy in stroboscopic time. Moreover, for the special example of a $sl(2)$ Floquet CFT, an OTOC computation in large-$c$ CFTs in the heating phase also shows a Lyapunov growth, characteristic of early time chaos Das:2022jrr . In the bulk description, such a Lyapunov growth in a holographic CFT is understood to be a consequence of blue-shifting of energy in the presence of a horizon Shenker:2013pqa . Indeed, as was confirmed in Das:2022pez , there is an observer-dependent horizon in the dual description of the $sl(2)$ Floquet CFT in the heating phase  [^1] For other works on the holographic bulk dual of Floquet CFTs or deformed CFTs see deBoer:2023lrd ; Jiang:2024hgt ; Kudler-Flam:2023ahk ; MacCormack:2018rwq .. As was shown in that paper, one can reproduce the Lyapunov exponent from a bulk computation, following the methods of Shenker:2013pqa .
@@ 2.2 Heating phase as Modular Hamiltonian
Since the quadratic Casimir is negative, we can conclude that the Hamiltonian in the heating phase corresponds to the extended Modular Hamiltonian of an interval in the vacuum state of a CFT, defined on a ring of length L. In Tada:2019rls , the authors considered a simpler choice where $\beta=\gamma$. In that case,
@@
Thus, the Floquet dynamics in the heating phase describes the evolution of operators in modular time111111now identified with the stroboscopic time, inside the corresponding Entanglement Wedge. This, Modular dynamics, can be better studied by going to a conformal frame ($CF_{2}$), where the modular time is identified as time. The conformal map to this new frame is given by:
@@
To describe a consistent theory on the sub-region, one needs to impose boundary conditions on the fixed points, after cutting off a small region ($\epsilon$) around itOhmori:2014eia . Imposing conformal boundary conditions, one ends up with a BCFT on the cylinder, with the Modular Hamiltonian generating translations along the Euclidean modular time circle 121212since the Modular Hamiltonian generates rotations around the fixed points Cardy:2016fqc )..
By definition, the BCFT on the Euclidean cylinder represents a thermal state, whose thermal entropy is the same as the entanglement entropy of the interval. This construction has been used to compute the entanglement entropy in 2D CFTs Holzhey:1994we . The spectrum of the Modular Hamiltonian would then be the thermal spectrum of a BCFT of length $l=2\log\left[\frac{L\sin\frac{\pi(R_{2}-R_{1})}{L}}{\pi\epsilon}\right]$. In the limit $\epsilon\rightarrow 0$, this becomes effectively an infinite length system and therefore one expects that this spectrum to become continuous in this limit  [^13] This provides us with an effective regulator for 2D CFTs on an infinite line Tada:2019rls . See SDasToAppear , for a detailed analysis of the spectrum in this Quantization..
@@ 3.1 AdS Rindler description of the heating phase
Therefore, in this case, the Schwarzians of the transformations are 00. As a result, after the analytic continuation of $\sigma^{0}$, the metric is:
@@ 4.1 A brief review of the Heating-Phase Quantization
Given a heating phase Hamiltonian (4) with $\delta<0$, one can use the prescription by Tada et al. Ishibashi:2016bey ; Tada:2019rls to define the conserved charges corresponding to new time translation generated by (4), as the following
@@
However, it is important to note that the function $f_{n}(z)$ depends crucially on the choice of $g(z)$. Therefore, unlike the usual Virasoro generators the label $n$ can take any value (continuous or discrete) depending on the final expression of $f_{n}(z)$ (for a given $g(z)$) and its single valued-ness on the complex plane. For the heating phase, Tada:2019rls showed that the $f_{n}(z)$ has divergences at the fixed points of the dynamics. This makes the modes $\mathcal{L}_{n}$ to be non-analytic at the fixed points. In particular, to obtain the algebra of $\mathcal{L}_{n}$s, one needs to take care of these divergences. Therefore, one needs to introduce a cut-off $\epsilon$ circling the fixed points. Thus, introduction of cut-off will modify the constant-time contour $C\rightarrow C_{\epsilon}$ and the eigenmodes $\mathcal{L}_{n}\rightarrow\mathcal{L}_{n}^{\epsilon}$. Once we demand the $\mathcal{L}_{n}^{\epsilon}$s satisfy Virasoro algebra with finite central extension term  [^19] The corresponding classical space-time generators will satisfy the Witt algebra Ishibashi:2016bey , the generator labels $n$ in (28) are restricted to take discrete values, rendering the spectrum discrete Tada:2019rls  [^20] See SDasToAppear for a review and the final form of the modified Virasoro algebra.. In particular, $n$ takes an integer or half-integer value multiplied by a $\epsilon$ dependent fractional term. After a proper rescaling Tada:2019rls , one would end up with the following form of the Virasoro algebra, where $n\in\mathbb{Z}$:
@@
It is noteworthy to mention that, due to the boundary condition placed at the cut-off, the whole problem is getting mapped to a BCFT on a finite cylinder as we mentioned earlier. As argued in SDasToAppear , the boundary conditions on those two cut-offs can be fixed by conformal boundary conditions. Hence, from the usual BCFT point of view, we will end up with a single copy of Virasoro algebra where the holomorphic and anti-holomorphic Virasoro generators would not be independent anymore. This makes (30) to be a unique and distinct Virasoro algebra from the usual one.
@@ 4.2 Bulk realization of the Virasoro algebra
5 Entanglement Entropy as BH Entropy
In this section, we will highlight that the entanglement entropy between the two fixed points is equal to the Bekenstein Hawking entropy corresponding to the AdS-Rindler metric. More specifically, we will compute the holographic Entanglement entropy by computing the length of the RT surface and comparing it with the Rindler entropy.
@@ 6 Discussion
In this article, we have explored some consequences that follow from the identification of the heating phase Hamiltonian with the Modular Hamiltonian of a sub-region, in the vacuum state. From this perspective, the heating phase quantization describes the evolution of operators inside the causal diamond of the subregion, between two fixed points of the $sl(2,\mathbb{R})$ flow, in modular/stroboscopic time. To study the modular quantization of this subregion, one needs to transform to a conformal frame, wherein the time coordinate is identified with the modular time. One then imposes a conformal boundary condition at an $\epsilon$ cut-off away from the fixed points of the Floquet Hamiltonian flows. This reduces the system of study to that of a BCFT on a strip. We also showed that holographically, this BCFT describes the dynamics inside the RT Entanglement Wedge, described by an AdS Rindler metric, where the Rindler Horizon is identified with the RT surface, and that the boundary cut-off, near the fixed points, translates to a cut-off, analogous to a stretched horizon [^22] See SDasToAppear and references therein where the connection to the stretched horizon has been explored in details., near the RT surface or the Rindler horizon. We further argued that the single copy of the conformal algebra preserved by the boundary conditions of the BCFT should be dual to the single copy of the Virasoro symmetry at the stretched Horizon. From the perspective of driven CFTs, we also argued that the different dynamical phases of these $CFT$s can be classified by the type of Von Neumann factors. In particular as one moves from the non-heating to the heating phase, by tuning the external control parameters like the frequency or time period of the drive, the factor changes from type $I$ to type $III$. This, again, is a direct consequence of the identification of the heating phase dynamics with Modular quantization.
@@
Since in the bulk description, the qualitative differences in the three dynamical phases correspond to differences in observations made by three different classes of observers Das:2022pez , it follows that the Von-Neumann algebra of observables accessible to the observer, dual to the heating phase would be type $III$, while the corresponding Von-Neumann algebra of observables accessible to the observer dual to the non-heating phase should be Type $I$. It would be interesting to understand this better along the lines of recent investigations Witten:2023qsv ; Penington:2023dql ; Strohmaier:2023opz ; Chen:2024rpx .
@@
Given that the boundary theory in our work describes the AdS Rindler Black hole, it should be possible to explicitly construct the microstates of the black hole directly from the heating phase of the driven CFT. In fact, these geometries would be the so-called fixed area states, since these would be eigenstates of the $H^{(u)}_{B_{1}B_{2}}$, which in the large $c$ is identified with the area operatorDong:2018seb . It would be nice to explicitly construct these states directly in the CFT [^23] Partial progress along this line has been done in SDasToAppear ..
@@
It would also be interesting to understand better the connection between the ‘emergent’ horizon Virasoro algebra and the ‘emergent’ boundary Virasoro algebra in Modular quantization. In fact, as shown in SDasToAppear , the microcanonical counting of states of this Hamiltonian which also describes the horizon CFT, would exactly reproduce the Bekenstein-Hawking entropy of BTZ. In the context of horizon CFT, this implies that these states live in the (vicinity of) horizon rather than at an asymptotic boundary. Also, a similar structure of near horizon Virasoro algebra exists in any higher dimensional asymptotic AdS or flat spacetime, as argued in Solodukhin:1998tc . Hence, we suspect the appearance of this conformal field theory and the emergence of horizon Virasoro symmetry have far-reaching consequences, beyond AdS/CFT. We hope to explore some of these issues in the immediate future.

### v2

@@ Abstract
In this article, we explore the connection between the heating phase of periodically driven CFTs and the Modular Hamiltonian of a subregion in the vacuum state. We show that the heating phase Hamiltonian corresponds to the Modular Hamiltonian, with the fixed points mapping to the endpoints of the subregion. In the bulk dual, we find that these fixed points correspond to the Ryu-Takayanagi surface of the AdS-Rindler wedge. Consequently, the entanglement entropy associated to the boundary interval within two fixed points exactly matches with the Rindler entropy of AdS-Rindler. We observe the emergent Virasoro algebra in the boundary quantization of the Modular Hamiltonian has a striking similarity with the emergent near Horizon Virasoro algebra. This is a consequence of the fact that while obtaining the boundary Virasoro algebra, a cut-off with conformal boundary condition around the fixed point is introduced, which in the bulk is related to a stretched horizon, with an emergent two-dimensional conformal symmetry. We also argue that as one tunes the parameter space of Floquet Hamiltonians to transition from the non-heating to the heating phase the operator algebra type changes from Von Neumann type $I$ to $III_{1}$ factor, providing a non-equilibrium analogue of the Hawking-Page transition.
@@ 1 Introduction
The study of Floquet CFTs has recently garnered considerable attention as analytically tractable models of physics out of equilibrium. At stroboscopic times, the evolution of such systems is governed by an effective Hamiltonian, referred to in the literature as the Floquet Hamiltonian $H_{F}$. One of the novel features of their dynamics is the existence of three different dynamical phases, characterized by qualitatively different temporal growths of the entanglement entropy and energy. In particular, one of the phases, aptly referred to as the heating phase, is characterized by an exponentially growing energy and linearly growing entanglement entropy in stroboscopic time. Moreover, for the special example of a $sl(2)$ Floquet CFT, an OTOC computation in large-$c$ CFTs in the heating phase also shows a Lyapunov growth, characteristic of early time chaos Das:2022jrr . In the bulk description, such a Lyapunov growth in a holographic CFT is understood to be a consequence of blue-shifting of energy in the presence of a horizon Shenker:2013pqa . Indeed, as was confirmed in Das:2022pez , there is an observer-dependent horizon in the dual description of the $sl(2)$ Floquet CFT in the heating phase  [^1] For other works on the holographic bulk dual of Floquet CFTs or deformed CFTs see deBoer:2023lrd ; Jiang:2024hgt ; Kudler-Flam:2023ahk ; MacCormack:2018rwq ; Caputa:2022zsr .. As was shown in that paper, one can reproduce the Lyapunov exponent from a bulk computation, following the methods of Shenker:2013pqa .
@@ 2.2 Heating phase as Modular Hamiltonian
Since the quadratic Casimir is negative, we can conclude that the Hamiltonian in the heating phase corresponds to the extended Modular Hamiltonian of an interval in the vacuum state of a CFT, defined on a ring of length L  [^11] This point was previously mentioned in Lapierre:2019rwj .. In Tada:2019rls , the authors considered a simpler choice where $\beta=\gamma$. In that case,
@@
Thus, the Floquet dynamics in the heating phase describes the evolution of operators in modular time121212now identified with the stroboscopic time, inside the corresponding Entanglement Wedge. This, Modular dynamics, can be better studied by going to a conformal frame ($CF_{2}$), where the modular time is identified as time. The conformal map to this new frame is given by:
@@
To describe a consistent theory on the sub-region, one needs to impose boundary conditions on the fixed points, after cutting off a small region ($\epsilon$) around itOhmori:2014eia . Imposing conformal boundary conditions, one ends up with a BCFT on the cylinder, with the Modular Hamiltonian generating translations along the Euclidean modular time circle 131313since the Modular Hamiltonian generates rotations around the fixed points Cardy:2016fqc )..
By definition, the BCFT on the Euclidean cylinder represents a thermal state, whose thermal entropy is the same as the entanglement entropy of the interval. This construction has been used to compute the entanglement entropy in 2D CFTs Holzhey:1994we . The spectrum of the Modular Hamiltonian would then be the thermal spectrum of a BCFT of length $l=2\log\left[\frac{L\sin\frac{\pi(R_{2}-R_{1})}{L}}{\pi\epsilon}\right]$. In the limit $\epsilon\rightarrow 0$, this becomes effectively an infinite length system and therefore one expects that this spectrum to become continuous in this limit  [^14] This provides us with an effective regulator for 2D CFTs on an infinite line Tada:2019rls . See Das:2024mlx , for a detailed analysis of the spectrum in this Quantization..
@@ 3.1 AdS Rindler description of the heating phase
Therefore, in this case, the Schwarzians of the transformations are zero. As a result, after the analytic continuation of $\sigma^{0}$, the metric is:
@@ 4.1 A brief review of the Heating-Phase Quantization
Given a heating phase Hamiltonian (4) with $\delta<0$, one can use the prescription in Ishibashi:2016bey ; Tada:2019rls to define the conserved charges corresponding to new time translation generated by (4), as the following
@@
However, it is important to note that the function $f_{n}(z)$ depends crucially on the choice of $g(z)$. Therefore, unlike the usual Virasoro generators the label $n$ can take any value (continuous or discrete) depending on the final expression of $f_{n}(z)$ (for a given $g(z)$) and its single valued-ness on the complex plane. For the heating phase, Tada:2019rls showed that the $f_{n}(z)$ has divergences at the fixed points of the dynamics. This makes the modes $\mathcal{L}_{n}$ to be non-analytic at the fixed points. In particular, to obtain the algebra of $\mathcal{L}_{n}$s, one needs to take care of these divergences. Therefore, one needs to introduce a cut-off $\epsilon$ circling the fixed points. Thus, introduction of cut-off will modify the constant-time contour $C\rightarrow C_{\epsilon}$ and the eigenmodes $\mathcal{L}_{n}\rightarrow\mathcal{L}_{n}^{\epsilon}$. Once we demand the $\mathcal{L}_{n}^{\epsilon}$s satisfy Virasoro algebra with finite central extension term  [^20] The corresponding classical space-time generators will satisfy the Witt algebra Ishibashi:2016bey , the generator labels $n$ in (28) are restricted to take discrete values, rendering the spectrum discrete Tada:2019rls  [^21] See Das:2024mlx for a review and the final form of the modified Virasoro algebra.. In particular, $n$ takes an integer or half-integer value multiplied by a $\epsilon$ dependent fractional term. After a proper rescaling Tada:2019rls , one would end up with the following form of the Virasoro algebra, where $n\in\mathbb{Z}$:
@@
It is noteworthy to mention that, due to the boundary condition placed at the cut-off, the whole problem is getting mapped to a BCFT on a finite cylinder as we mentioned earlier. As argued in Das:2024mlx , the boundary conditions on those two cut-offs can be fixed by conformal boundary conditions. Hence, from the usual BCFT point of view, we will end up with a single copy of Virasoro algebra where the holomorphic and anti-holomorphic Virasoro generators would not be independent anymore. This makes (30) to be a unique and distinct Virasoro algebra from the usual one.
@@ 4.2 Bulk realization of the Virasoro algebra
5 Entanglement Entropy as AdS-Rindler Entropy
In this section, we will highlight that the entanglement entropy between the two fixed points is equal to the Rindler entropy corresponding to the AdS-Rindler metric. More specifically, we will compute the holographic Entanglement entropy by computing the length of the RT surface and comparing it with the Rindler entropy.
@@ 6 Discussion
In this article, we have explored some consequences that follow from the identification of the heating phase Hamiltonian with the Modular Hamiltonian of a sub-region, in the vacuum state. From this perspective, the heating phase quantization describes the evolution of operators inside the causal diamond of the subregion, between two fixed points of the $sl(2,\mathbb{R})$ flow, in modular/stroboscopic time. To study the modular quantization of this subregion, one needs to transform to a conformal frame, wherein the time coordinate is identified with the modular time. One then imposes a conformal boundary condition at an $\epsilon$ cut-off away from the fixed points of the Floquet Hamiltonian flows. This reduces the system of study to that of a BCFT on a strip. We also showed that holographically, this BCFT describes the dynamics inside the RT Entanglement Wedge, described by an AdS Rindler metric, where the Rindler Horizon is identified with the RT surface, and that the boundary cut-off, near the fixed points, translates to a cut-off, analogous to a stretched horizon [^23] See Das:2024mlx and references therein where the connection to the stretched horizon has been explored in details., near the RT surface or the Rindler horizon. We further argued that the single copy of the conformal algebra preserved by the boundary conditions of the BCFT should be dual to the single copy of the Virasoro symmetry at the stretched Horizon. From the perspective of driven CFTs, we also argued that the different dynamical phases of these $CFT$s can be classified by the type of Von Neumann factors. In particular as one moves from the non-heating to the heating phase, by tuning the external control parameters like the frequency or time period of the drive, the factor changes from type $I$ to type $III$. This, again, is a direct consequence of the identification of the heating phase dynamics with Modular quantization.
@@
Since in the bulk description, the qualitative differences in the three dynamical phases correspond to differences in observations made by three different classes of observers Das:2022pez , it follows that the Von-Neumann algebra of observables accessible to the observer, dual to the heating phase would be type $III$, while the corresponding Von-Neumann algebra of observables accessible to the observer dual to the non-heating phase should be Type $I$. It would be interesting to understand this better along the lines of recent investigations Witten:2023qsv ; Penington:2023dql ; Strohmaier:2023opz ; Kudler-Flam:2024psh ; Chen:2024rpx .
@@
Given that the boundary theory in our work describes the AdS Rindler Black hole, it should be possible to explicitly construct the microstates of the black hole directly from the heating phase of the driven CFT. In fact, these geometries would be the so-called fixed area states, since these would be eigenstates of the $H^{(u)}_{B_{1}B_{2}}$, which in the large $c$ is identified with the area operatorDong:2018seb . It would be nice to explicitly construct these states directly in the CFT [^24] Partial progress along this line has been done in Das:2024mlx ..
@@
It would also be interesting to understand better the connection between the ‘emergent’ horizon Virasoro algebra and the ‘emergent’ boundary Virasoro algebra in Modular quantization. In fact, as shown in Das:2024mlx , the microcanonical counting of states of this Hamiltonian which also describes the horizon CFT, would exactly reproduce the Bekenstein-Hawking entropy of BTZ. In the context of horizon CFT, this implies that these states live in the (vicinity of) horizon rather than at an asymptotic boundary. Also, a similar structure of near horizon Virasoro algebra exists in any higher dimensional asymptotic AdS or flat spacetime, as argued in Solodukhin:1998tc . Hence, we suspect the appearance of this conformal field theory and the emergence of horizon Virasoro symmetry have far-reaching consequences, beyond AdS/CFT. We hope to explore some of these issues in the immediate future.